OpenStudy (anonymous):

What is 0.56 written as a fraction in simplest form? A. 2/5 B. 14/25 C. 1/2 D. 5 3/5

OpenStudy (shadowlegendx):

What do you think the answer is?

OpenStudy (anonymous):

Well 0.56 is the same as 56/100

OpenStudy (anonymous):

but it's not in the choises

OpenStudy (shadowlegendx):

Yes

OpenStudy (anonymous):

now that I think....it's B right?

OpenStudy (anonymous):

@ShadowLegendX

OpenStudy (shadowlegendx):

Ok so I'm going to explain the process

OpenStudy (shadowlegendx):

What you need to do here is reducing the fractions

OpenStudy (paki):

yeah 56/100 = 0.56 = 14/25 = 0.56....

OpenStudy (shadowlegendx):

The way this works is say you have 5/10 it is the same as 1/2 

OpenStudy (anonymous):

I knew it!

OpenStudy (shadowlegendx):

What you would have done is reducing it by doing \[56 \div 4 = 14\] \[100 \div 4 = \]

OpenStudy (shadowlegendx):

25

OpenStudy (shadowlegendx):

Then we would have 14/25